Management Accounting

5. Peach Cos latest results are as follows: Profit before interest and taxation Profit before taxation Profit after tax Rs.
Answer #1

Answer is option c . 20 %

Return on capital employed formula is calculated by dividing net operating profit or EBIT by the employed capital.

Capital employed is derived by subtracting current liabilities from total assets; or alternatively by adding noncurrent liabilities to owners' equity

Capital employed 10000 + 2500 = 12500$ ( non current liab + equity)

EBIT. 2500$

=2500/12500. = 20%